Crane Company gathered the following data about the three products that it produces: Product Present Sales Value Estimated Additional Processing Costs Estimated Sales if Processed Further A $12800 $8400 $22200 B 14800 5400 19200 C 11800 3400 17200 Which of the products should not be processed further? Product C Product A Product B Products A and C
Answer #1

Incremental revenue

Product A = (22,200-8,400) - 12,800 = +1000

Product B = (19,200-5,400) - 14,500 = - 700

Product C = (17,200-3,400) - 11,800 = +2000

Product B shouldn't be processed further

Option C (Product B)